# FastVID: Dynamic Density Pruning for Fast Video Large Language Models

### v1

@@ Abstract
Video Large Language Models have shown impressive capabilities in video comprehension, yet their practical deployment is hindered by substantial inference costs caused by redundant video tokens. Existing pruning techniques fail to fully exploit the spatiotemporal redundancy inherent in video data. To bridge this gap, we perform a systematic analysis of video redundancy from two perspectives: temporal context and visual context. Leveraging this insight, we propose Dynamic Density Pruning for Fast Video LLMs termed FastVID. Specifically, FastVID dynamically partitions videos into temporally ordered segments to preserve temporal structure and applies a density-based token pruning strategy to maintain essential visual information. Our method significantly reduces computational overhead while maintaining temporal and visual integrity. Extensive evaluations show that FastVID achieves state-of-the-art performance across various short- and long-video benchmarks on leading Video LLMs, including LLaVA-OneVision and LLaVA-Video. Notably, FastVID effectively prunes $\textbf{90}\%$ of video tokens while retaining $\textbf{98.0}\%$ of LLaVA-OneVision’s original performance. The code is available at https://github.com/LunarShen/FastVID.
@@ 1 Introduction
Video Large Language Models (Video LLMs) [5, 17, 14, 43, 35] have gained prominence for video comprehension. These models process videos by encoding sampled frames into video tokens through a vision encoder. These tokens are typically pooled before being integrated with query tokens and passed to an LLM for response generation. Despite their effectiveness, the high computational cost of inference remains a significant challenge. While some methods [29, 22, 37, 13, 28, 45] introduce trainable compression modules to mitigate this issue, they require extensive retraining, leading to high training costs. In contrast, we focus on an inference-time acceleration strategy that enhances efficiency without requiring additional training.
@@
Figure 1: Effective token compression in Video LLMs relies on preserving both temporal and visual context. (a) shows the impact of temporal disruptions on video understanding, emphasizing the importance of temporal structure preservation. (b–d) show the effects of different token selection strategies for visual merging (Patches that share the same inner and border color are merged.), indicating that Density-based Sampling retains both distinctive and representative context while effectively reducing redundancy.
@@
This computational burden is primarily caused by the high volume of video tokens, making effective token compression essential. While prior image compression methods [4, 26, 42, 38, 3] reduce redundancy within a single image, they fail to exploit temporal dependencies across frames. Meanwhile, video compression methods like DyCoke [31] and PruneVid [11] primarily prune at the pre-filling and decoding stages, treating video tokens like textual tokens and neglecting video-specific characteristics. There remains a notable gap in the systematic exploration of video redundancy itself. A more video-aware pruning approach is necessary to explicitly exploit spatiotemporal redundancy while preserving essential semantic content.
To address this limitation, we focus on minimizing spatiotemporal redundancy at the input stage. In particular, we systematically analyze video redundancy from two key perspectives: temporal context and visual context.
@@
Temporal context is fundamental to video understanding, as the order and continuity of frames directly influence semantic interpretation. As depicted in Figure 1a, disrupting frame order (shuffled) or omitting frames (incomplete) leads to incorrect comprehension, highlighting the necessity of preserving temporal structure. To achieve this, we segment the video into temporally ordered segments, grouping highly similar consecutive frames. Pruning is applied within each high-redundancy segment but not across segments, preserving the essential temporal structure for accurate video understanding.
@@
Within each segment, it is crucial to merge redundant tokens while preserving visual context. A state-of-the-art method VisionZip [38] applies uniform token sampling followed by merging redundant tokens [3] as shown in Figure 1c. However, this strategy is highly sensitive to token selection, potentially leading to the loss of important details, such as the guitar body being incorrectly merged with the background. To address this, we propose a density-based token sampling in Figure 1d. Specifically, high-density tokens, surrounded by numerous similar tokens, serve as candidates for selection. We select density peak tokens as anchors, which have a higher density than their neighbors and by a relatively large distance from tokens with higher densities [24]. This strategy ensures that the selected tokens are both representative and distinctive, effectively preserving segment visual context while reducing redundancy.
@@
Building on these insights, we propose Dynamic Density Pruning for Fast Video LLMs (FastVID), which optimizes video pruning by dynamically segmenting frames and applying density-based pruning within each segment. We begin with Dynamic Temporal Segmentation, which adaptively partitions videos into segments. Within each segment, we introduce Density Spatiotemporal Pruning to retain both global visual context and salient details. FastVID significantly accelerates Video LLM inference while preserving both temporal and visual integrity.
To assess the generalization capability of our approach, we evaluate it on LLaVA-OneVision [14] and LLaVA-Video [43], two leading Video LLMs. Notably, LLaVA-Video11footnotemark: 1 employs newline tokens based on the positional information of video tokens, which is an aspect that many pruning methods do not explicitly consider. Our method effectively adapts to this design, making it particularly well-suited for LLaVA-Video. To evaluate the effectiveness of our method, we perform extensive experiments on MVBench [16], LongVideoBench [36], MLVU [44], and VideoMME [10]. These benchmarks cover a wide range of video complexities and durations, ensuring a comprehensive evaluation. Notably, FastVID prunes 90% of video tokens while maintaining 98.0% of LLaVA-OneVision’s original performance across all benchmarks.
The main contributions are summarized as follows:
  • We analyze Video LLM compression from both temporal and visual context perspectives, emphasizing the importance of maintaining temporal and visual integrity.
@@
  • We propose FastVID, a novel pruning framework that employs Dynamic Temporal Segmentation to partition videos into temporally ordered segments and Density Spatiotemporal Pruning to retain global segment information and key details.
  • Our FastVID generalizes across different Video LLM architectures and achieves state-of-the-art performance on diverse video benchmarks.
@@ 2 Related Work
Video LLMs. With the rapid advancement of LMMs [1, 6, 32] and MLLMs [33, 19, 20, 15, 2, 21], there has been growing interest in Video LLMs. Existing Video LLMs can be broadly categorized based on how they process video tokens: general Video LLMs and Video LLMs with training-time compression.
General Video LLMs [18, 5, 17, 14, 43, 35] directly process raw video tokens or apply pooling. Video-LLaVA [17] leverages shared projection layers to obtain unified visual representations. LLaVA-OneVision [14] demonstrates strong video understanding through task transfer from images. LLaVA-Video [43] creates a high-quality synthetic dataset for video instruction-following. To better capture the spatiotemporal structure of video, some models [43, 35] introduce additional designs for video positional information. Qwen2-VL [35] propose M-RoPE to decompose rotary embedding into temporal, height, and width components. LLaVA-Video [^1] https://github.com/LLaVA-VL/LLaVA-NeXT introduces newline tokens to distinguish spatial and temporal positions effectively.
Video LLMs with training-time compression [29, 22, 37, 13, 28, 45] aim to significantly reduce the number of video tokens, enabling longer video sequences. PLLaVA [37] introduces an adaptive average structure pooling to extend image LLMs. Video-ChatGPT [22] extracts both spatial and temporal features through temporal and spatial pooling respectively. Chat-UniVi [13] progressively clusters visual tokens and provides multi-scale features. MovieChat [29] introduces a memory management mechanism to enhance long video understanding. LongVU [28] employs cross-modal query and inter-frame dependencies to adaptively reduce video redundancy. Apollo [45] explores scaling consistency and uses the Perceiver Resampler [12].
Figure 2: The overview of our FastVID. DySeg dynamically partitions video tokens into temporally ordered, high-redundancy segments, while STPrune performs density-based pruning within each segment, enhancing efficiency without loss of key information.
However, general Video LLMs remain the dominant paradigm, with LLaVA-OneVision being widely adopted due to its adaptability and superior performance. Moreover, Video LLMs with training-time compression often exhibit lower performance compared to their general counterparts. For example, Apollo-7B with trainable compression achieves $61.3$ on VideoMME (wo sub.), whereas LLaVA-Video-7B achieves $63.3$.
Therefore, we focus on inference-time acceleration for general Video LLMs, aiming to improve computational efficiency while maintaining strong performance. To evaluate our method, we select LLaVA-OneVision and LLaVA-Video as representative models.
@@
Recent studies [4, 26, 42, 38, 3] primarily focus on spatial token reduction to accelerate Image LLMs. FastV [4] selects text-relevant tokens at shallow layers of the LLM. LLaVA-PruMerge [26] uses attention scores from the [CLS] token to prune spatial redundancy. SparseVLM [42] proposes a token recycling strategy to aggregate and reconstruct tokens to be pruned. VisionZip [38] reduces visual redundancy in the vision encoders. However, these methods overlook the temporal relationships across frames.
Due to the high volume of video tokens in Video LLMs, recent video compression methods [31, 11] have gained increasing attention. DyCoke [31] merges tokens across frames and applies dynamic KV cache reduction. PruneVID [11] clusters video tokens and selects those most relevant to query tokens. However, pruning in these methods predominantly occurs during the prefilling and decoding stages, where spatiotemporal characteristics of video data are not adequately considered.
In this paper, we focus on temporal and visual context for spatiotemporal token pruning at the input stage. By reducing video tokens before LLM processing, our FastVID significantly enhances computational efficiency while facilitating easy deployment, including compatibility with FlashAttention [7], KV cache, and a plug-and-play design for seamless integration into existing Video LLMs.
@@ 3.1 FastVID
To address these limitations, we propose FastVID, a novel method that accelerates inference in Video LLMs by reducing redundant tokens while preserving critical temporal and visual context. First, we introduce Dynamic Temporal Segmentation (DySeg) in Section 3.2, which partitions a video into temporally ordered segments with high redundancy. Within each segment, we introduce Density Spatiotemporal Pruning (STPrune) in Section 3.3, which retains essential visual context and salient details.
Figure 2 presents an overview of FastVID. Given an input video, $F$ frames are uniformly sampled (e.g., 32 in LLaVA-OneVision, 64 in LLaVA-Video). Each frame is individually processed by the vision encoder. The extracted tokens are projected and pooled into a video token sequence. With FastVID, we effectively eliminate redundant tokens while preserving critical information. The retained video sequence, combined with query tokens, is then fed into the LLM to generate the response.
Figure 3: Visualization of different segmentation methods on 12 sampled frames from a video example. Each frame’s top-left corner displays its sequential index. (b) Fixed-interval Segmentation struggles to maintain high intra-segment similarity, leading to visually diverse frames within the same segment. (c) Cluster-based Segmentation disrupts temporal order, grouping frames from different time periods to the same segment. (d) Our DySeg adaptively partitions the video, preserving both temporal structure and high intra-segment similarity. Please refer to Table  2 for a detailed quantitative comparison.
@@ 3.2 Dynamic Temporal Segmentation
Section 1 highlights the importance of video segmentation in maintaining temporal order across segments while ensuring high similarity within each segment.
As shown in Figure 3b and 3c, there are two common segmentation approaches for sampled frame sequences. In Figure 3b, Fixed-interval Segmentation partitions the sequence into segments of a fixed length, preserving temporal order but potentially grouping visually dissimilar frames. Figure 3c shows Cluster-based Segmentation, where frames are grouped into three clusters based on frame similarity. However, this approach suffers from a predefined cluster number, leading to ineffective segmentation when video complexity varies. As a result, the first segment contains similar objects but different scenes. Furthermore, clustering may disrupt the temporal order by ignoring critical temporal information, such as omitting key frames (e.g., the ninth frame) and incorrectly grouping frames from different time periods into the first segment.
To address these limitations, we propose Dynamic Temporal Segmentation, a simple yet effective method that adaptively refines segment boundaries according to video complexity. DySeg achieves both temporal structure and high intra-segment similarity, generating fewer partitions for simple scenes and finer ones for more complex scenes.
To maintain high similarity within each segment, we aim to minimize similarity between adjacent segments. Thus, we segment the video based on transition similarity between adjacent frames. Specifically, we utilize global frame features $\mathbf{f}$ to compute the cosine similarity:
$t_{i}=\cos(\mathbf{f}_{i},\mathbf{f}_{i+1}),\quad i=1,\dots,F-1,$
$\mathbf{T}=\{t_{1},t_{2},\dots,t_{F-1}\},$ (1)
where $t_{i}$ denotes the transition similarity between the $i$-th and $(i+1)$-th frame. $\mathbf{T}$ denotes the $F-1$ transition similarities for a sequence of $F$ frames. To achieve fine-grained segmentation, we select transitions that satisfy the following conditions:
@@
where $c$ denotes the minimum number of segments and $\tau$ denotes the threshold for transition similarity. $\mathbf{S}_{1}$ denotes the $c-1$ least similar frame transitions, while $\mathbf{S}_{2}$ denotes transitions where similarity falls below a fixed threshold $\tau$. The final segmentation is determined by the union $\mathbf{S}$. In simple videos with minimal shot transitions, $\mathbf{S}_{2}$ is often empty. In such cases, $\mathbf{S}_{1}$ enables finer segmentation by distinguishing subtle temporal changes. For complex videos, $\mathbf{S}_{1}$ is typically a subset of $\mathbf{S}_{2}$, where $\mathbf{S}_{2}$ ensures that adjacent frames with similarity below $\tau$ are assigned to different segments. In Figure 3d, our DySeg effectively segments videos by dynamically adjusting granularity in a simple yet effective manner, outperforming Fixed-interval and Clustering-based methods.
@@ 3.3 Density Spatiotemporal Pruning
After obtaining segments with highly similar frames, we introduce STPrune to reduce redundant tokens. It consists of two key modules: Density-based Token Merging (DTM) for the segment’s visual context and Attention-based Token Selection (ATS) for salient details. In each segment, we retain $R\%$ of tokens, where $dR\%$ are obtained using DTM and $(1-d)R\%$ are obtained using ATS. $d$ controls the token allocation between the two modules.
Density-based Token Merging. For intra-segment merging, we first identify anchor tokens and to-be-merged tokens. Selecting anchor tokens from the entire segment would disrupt their spatial relationships, as anchor tokens corresponding to different components of an object might be distributed across multiple frames. To mitigate this, we restrict anchor token selection to specific frames. Specifically, we sample anchor frames at a fixed interval $p$ and select anchor tokens from these frames. The remaining tokens in the segment are treated as to-be-merged tokens. For a segment of $f$ frames, the total number of anchor frames is $\lfloor f/p\rfloor$. From each anchor frame, $dR/\lfloor f/p\rfloor\%$ of tokens are selected as anchor tokens.
@@
Figure 4b illustrates DTM when the segment length is 1 and shows that it consists of three key steps. First, we follow density peaks clustering algorithms [24, 9] to compute the density score. For each token in the anchor frame $[v_{1},v_{2},\cdots,v_{N}]$, we calculate its local density $\rho_{i}$ and its distance to the closest higher-density token $\delta_{i}$, obtaining the final density score $\rho_{i}\times\delta_{i}$:
$\rho_{i}=\exp(-\frac{1}{k}\sum_{v_{j}\in\text{kNN}(v_{i})}\text{d}(v_{i},v_{j})^{2})$ (3)
$\delta_{i}=\begin{cases}\min\limits_{j:\rho_{i}>\rho_{j}}~{}\text{d}(v_{i},v_{j}),&\text{if }\exists j\text{ s.t. }\rho_{i}>\rho_{j}\\ ~{}~{}\max\limits_{j}~{}~{}\text{d}(v_{i},v_{j}),&\text{otherwise}\end{cases}$ (4)
where $\text{d}(v_{i},v_{j})$ denotes the Euclidean distance. Density peak tokens with high $\rho_{i}\times\delta_{i}$ serve as anchor tokens, indicating that they are surrounded by neighbors with lower local density while being relatively distant from other high-density tokens. This ensures anchor tokens are both representative and distinctive.
Next, Similarity-based Assignment assigns each remaining token to its nearest anchor. Finally, we apply Anchor-centric Aggregation to merge the assigned tokens into their respective anchors, preserving key visual structures through representative tokens. For an anchor $a$ and its associated tokens $[b_{1},\cdots,b_{n}]$, the updated $a^{*}$ is computed as:
$a^{*}=\beta a+\frac{1-\beta}{n}\sum_{i=1}^{n}b_{i},$ (5)
where $\beta$ controls the balance between the anchor token and its associated tokens. In Figure 4, we compare our DTM with Cluster-based Token Merging. In LLMs, RoPE [30] encodes relative positional relationships between tokens, making positional information essential for maintaining the spatiotemporal structure of video tokens. While prior methods [11] rely on Cluster-based Token Merging, they discard positional information, causing RoPE to struggle with encoding spatiotemporal structure. In contrast, our DTM maintains the positional information of merged tokens, enhancing visual context understanding. This is particularly beneficial for LLaVA-Video, which explicitly depends on positional information, where our DTM significantly outperforms other methods (see Table 4).
@@
Specifically, for a given segment, we extract the [CLS] attention score from the pretrained SigLIP head, $\mathbf{A}\in\mathbb{R}^{f\times H\times W}$, where $f$ is the number of frames, and $H$ and $W$ are the spatial dimensions of frame tokens. Since Video LLMs incorporate pooling (see Figure 2), we apply the same operation to $\mathbf{A}$, resulting in $\bar{\mathbf{A}}\in\mathbb{R}^{f\times\bar{H}\times\bar{W}}$, where $\bar{H}$ and $\bar{W}$ are the pooled spatial dimensions. Finally, for each frame, we select the top $(1-d)R\%$ of tokens with the highest attention scores to preserve critical visual details.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Video Large Language Models (Video LLMs) [6, 19, 16, 47, 37] have shown strong performance in video understanding but incur substantial inference costs. While several methods [31, 24, 40, 15, 30, 49] explore training-time compression to mitigate this issue, they often require expensive retraining. In this work, we focus on an inference-time acceleration strategy that enhances efficiency without requiring additional training.
@@
Figure 1: Effective token compression in Video LLMs relies on preserving both temporal and visual context. (a) shows the impact of temporal disruptions, emphasizing the importance of temporal structure preservation. (b) show the effects of different token selection strategies for visual merging, where patches that share the same inner and border color are merged. Density-based Sampling retains both distinctive and representative context while effectively reducing redundancy.
@@
This computational burden is primarily caused by the high volume of video tokens, making effective token compression essential. While prior image compression methods [5, 28, 46, 42, 4] reduce redundancy within a single image, they fail to exploit temporal dependencies across frames. As a result, the nature of video redundancy remains underexplored. In this work, we systematically analyze video redundancy from two key perspectives: temporal context and visual context.
@@
Temporal context is fundamental to video understanding, as the order and continuity of frames directly influence semantic interpretation. As depicted in Figure 1(a), disrupting frame order (shuffled) or omitting frames (incomplete) leads to incorrect comprehension, highlighting the necessity of preserving temporal structure. To achieve this, we segment the video into temporally ordered segments, grouping highly similar consecutive frames. Pruning is applied within each high-redundancy segment but not across segments, preserving the essential temporal structure for accurate video understanding.
@@
To maintain visual context, redundant tokens within each segment must be effectively merged. A state-of-the-art method VisionZip [42] applies uniform token sampling followed by merging redundant tokens [4] as shown in Figure 1(b1). However, its token selection strategy is content-agnostic, potentially leading to the loss of important details, such as the guitar body being incorrectly merged with the background. To address this, we propose a density-based token sampling in Figure 1(b2). Specifically, high-density tokens, surrounded by numerous similar tokens, serve as candidates for selection. We select density peak tokens as anchors, which have a higher density than their neighbors and by a relatively large distance from tokens with higher densities [26]. This strategy ensures that the selected tokens are both representative and distinctive, effectively preserving segment visual context while reducing redundancy.
@@
Building upon these insights, we propose Dynamic Density Pruning for Fast VIDeo LLMs, minimizing spatiotemporal redundancy while preserving essential semantics. We begin with Dynamic Temporal Segmentation, which adaptively partitions videos into segments. Within each segment, we introduce Density Spatiotemporal Pruning to retain both global visual context and salient details. FastVID significantly accelerates inference while preserving both temporal and visual integrity.
[…]
@@
The main contributions are summarized as follows: (1) We analyze Video LLM compression from both temporal and visual context perspectives, emphasizing the importance of maintaining temporal and visual integrity. (2) We propose FastVID, a novel pruning framework that employs Dynamic Temporal Segmentation to partition videos into temporally ordered segments and Density Spatiotemporal Pruning to retain global segment information and key details. (3) Our FastVID achieves state-of-the-art performance across diverse video benchmarks and maintains robust accuracy even under extreme compression.
@@ 2 Related Work
Video LLMs. With the rapid advancement of LMMs [1, 7, 34] and MLLMs [35, 21, 22, 17, 3, 23], there has been growing interest in Video LLMs. These models can be categorized based on how they process video tokens: general Video LLMs and Video LLMs with training-time compression.
General Video LLMs [20, 6, 19, 16, 47, 37] directly process raw video tokens or apply pooling. Video-LLaVA [19] leverages shared projection layers to obtain unified visual representations. LLaVA-OneVision [16] demonstrates strong video understanding through task transfer from images. LLaVA-Video [47] creates a high-quality synthetic dataset for video instruction-following. To better capture the spatiotemporal structure of video, some models [47, 37] introduce additional designs for video positional information. LLaVA-Video [^1] https://github.com/LLaVA-VL/LLaVA-NeXT introduces newline tokens to distinguish spatial and temporal positions effectively.
Video LLMs with training-time compression [31, 24, 40, 15, 30, 49] aim to significantly reduce the number of video tokens, enabling longer video sequences. PLLaVA [40] introduces an adaptive average structure pooling to extend image LLMs. Video-ChatGPT [24] extracts both spatial and temporal features through temporal and spatial pooling respectively. Chat-UniVi [15] progressively clusters visual tokens and provides multi-scale features. LongVU [30] employs cross-modal query and inter-frame dependencies to adaptively reduce video redundancy. Apollo [49] explores scaling consistency and uses the Perceiver Resampler [14].
However, general Video LLMs remain the dominant paradigm, with LLaVA-OneVision being widely adopted due to its adaptability and superior performance. Therefore, we focus on inference-time acceleration for general Video LLMs. To evaluate our method, we select LLaVA-OneVision and LLaVA-Video as representative models.
@@
Recent studies [5, 28, 46, 42, 4, 39] primarily focus on spatial token reduction to accelerate Image LLMs. FastV [5] selects text-relevant tokens at shallow layers of the LLM. LLaVA-PruMerge [28] uses attention scores from the [CLS] token to prune spatial redundancy. SparseVLM [46] proposes a token recycling strategy to aggregate and reconstruct tokens to be pruned. VisionZip [42] reduces visual redundancy in the vision encoders. However, these methods overlook the temporal relationships across frames.
Due to the high volume of video tokens in Video LLMs, recent video compression methods [12, 33, 13] have gained increasing attention. FrameFusion [12] applies both merging and pruning across successive shallow LLM layers, but repeated pruning operations adversely affect overall efficiency. DyCoke [33] merges tokens across frames and applies dynamic KV cache reduction. However, its pruning during the prefilling stage struggles to achieve substantial token reduction while maintaining accuracy. PruneVID [13] clusters video tokens and selects those most relevant to query tokens, but this dependency on clustering introduces significant latency during compression. Despite these advances, efficient and accurate pruning under large token reduction remains unsolved.
In this paper, we focus on preserving temporal and visual context for spatiotemporal token pruning at the prefilling stage. By reducing video tokens before LLM processing, our FastVID significantly enhances computational efficiency while facilitating easy deployment, including compatibility with FlashAttention [8], KV cache, and a plug-and-play design for seamless integration into existing Video LLMs. Our method achieves robust performance even under extreme compression rates, offering a practical solution for fast Video LLMs.
@@ 3.1 FastVID
To address these limitations, we propose FastVID, a novel method that accelerates inference by reducing redundant tokens while preserving critical temporal and visual context. First, we introduce Dynamic Temporal Segmentation (DySeg) in Section 3.2, which partitions a video into temporally ordered segments with high redundancy. Within each segment, we introduce Density Spatiotemporal Pruning (STPrune) in Section 3.3, which retains essential visual context and salient details.
Figure 2: The overview of our FastVID.
Figure 2 presents an overview of FastVID. Given an input video, $F$ frames are uniformly sampled (e.g., 32 in LLaVA-OneVision, 64 in LLaVA-Video). Each frame is individually processed by the vision encoder. The extracted tokens are projected and pooled into a video token sequence. With FastVID, we effectively eliminate redundant tokens while preserving critical information. Specifically, DySeg dynamically partitions video tokens into temporally ordered, high-redundancy segments, while STPrune performs density-based pruning within each segment, enhancing efficiency without loss of key information. The final retained video tokens, combined with query tokens, is then fed into the LLM to generate the response.
@@ 3.2 Dynamic Temporal Segmentation
As shown in Figure 3(b) and 3(c), there are two common static segmentation methods for sampled frame sequences. In Figure 3(b), Fixed-interval Segmentation partitions the sequence into segments of a fixed length, preserving temporal order but potentially grouping visually dissimilar frames. Figure 3(c) shows Cluster-based Segmentation, where frames are grouped into three clusters based on frame similarity. However, it suffers from a predefined cluster number, leading to ineffective segmentation when video complexity varies. As a result, the first segment contains similar objects but different scenes. Furthermore, clustering may disrupt the temporal order by ignoring critical temporal information, such as omitting key frames (e.g., the 9th frame) and incorrectly grouping frames from different time periods into the first segment.
To address the limitations of static methods, we propose Dynamic Temporal Segmentation, a simple yet effective method that adaptively refines segment boundaries according to video complexity. DySeg achieves both temporal structure and high intra-segment similarity, generating fewer partitions for simple scenes and finer ones for more complex scenes.
To enable effective pruning, DySeg promotes high spatiotemporal redundancy within each segment by minimizing similarity between adjacent segments. Thus, we segment the video based on transition similarity between adjacent frames. Specifically, we utilize global frame features $\mathbf{f}$ to compute the cosine similarity:
$t_{i}=\cos(\mathbf{f}_{i},\mathbf{f}_{i+1}),\quad i=1,\cdots,F-1,$
$\mathbf{T}=\{t_{1},t_{2},\cdots,t_{F-1}\},$ (1)
where $t_{i}$ denotes the transition similarity between the $i$-th and $(i+1)$-th frame. $\mathbf{T}$ denotes the $F-1$ transition similarities for $F$ sampled frames. To achieve fine-grained segmentation, we select transitions that satisfy the following conditions:
@@
where $c$ denotes the minimum number of segments and $\tau$ denotes the threshold for transition similarity. $\mathbf{S}_{1}$ denotes the $c-1$ least similar frame transitions, while $\mathbf{S}_{2}$ denotes transitions where similarity falls below a fixed threshold $\tau$. Each transition in the union $\mathbf{S}$ marks a boundary between segments. In simple videos with minimal shot transitions, $\mathbf{S}_{2}$ is often empty. In such cases, $\mathbf{S}_{1}$ enables finer segmentation by distinguishing subtle temporal changes. For complex videos, $\mathbf{S}_{1}$ is typically a subset of $\mathbf{S}_{2}$, where $\mathbf{S}_{2}$ ensures that adjacent frames with similarity below $\tau$ are assigned to different segments. In Figure 3(d), our DySeg effectively segments videos by dynamically adjusting granularity in a simple yet effective manner, outperforming Fixed-interval and Clustering-based methods.
@@ 3.3 Density Spatiotemporal Pruning
After obtaining segments with highly similar frames, we introduce STPrune to reduce redundant tokens. It consists of two key modules: Density-based Token Merging (DTM) for segment visual context and Attention-based Token Selection (ATS) for salient details. For a segment of $P$ frames, we retain $rPN$ tokens in total, where $N$ is the number of tokens per frame, and $r$ denotes the retention ratio. These tokens are generated by DTM and ATS. Specifically, $drFN$ tokens are generated by DTM and $(1-d)rFN$ tokens are generated by ATS, where $d$ controls the token allocation between the two modules.
Density-based Token Merging. For segment visual context, we first identify anchor tokens and to-be-merged tokens. Selecting anchor tokens from the entire segment would disrupt their spatial relationships, as anchor tokens corresponding to different components of the same object might be distributed across multiple frames. To mitigate this, we restrict anchor token selection to specific frames. Specifically, we sample anchor frames at a fixed interval $p$ and select anchor tokens from these frames. The remaining tokens in the segment are treated as to-be-merged tokens. Specifically, the number of anchor frames is $\lceil P/p\rceil$. From each anchor frame, $drPN/\lceil P/p\rceil$ tokens are selected as anchor tokens. Please see Figure 5 for a visualization of DTM applied to video segments.
Figure 3: Visualization of different segmentation methods on 12 sampled frames from a video example. (b) Fixed-interval Segmentation struggles to maintain high intra-segment similarity, leading to visually diverse frames within the same segment. (c) Cluster-based Segmentation disrupts temporal order, grouping frames from different time periods to the same segment. (d) Our DySeg adaptively partitions the video, preserving both temporal structure and high intra-segment similarity. Please refer to Table 6 for a detailed quantitative comparison.
To further clarify DTM, Figure 4(b) illustrates DTM when the segment length is 1 and shows that it consists of three key steps. First, we follow density peaks clustering algorithms [26, 10] to compute the density score. For each token in the anchor frame $[v_{1},v_{2},\cdots,v_{N}]$, we calculate its local density $\rho_{i}$ and its distance to the closest higher-density token $\delta_{i}$, obtaining the final density score $\rho_{i}\times\delta_{i}$:
$\rho_{i}=\exp(-\frac{1}{k}\sum_{v_{j}\in\text{kNN}(v_{i})}\text{d}(v_{i},v_{j})^{2}),$ (3)
$\delta_{i}=\begin{cases}\min\limits_{j:\rho_{j}>\rho_{i}}~{}\text{d}(v_{i},v_{j}),&\text{if }\exists j\text{ s.t. }\rho_{j}>\rho_{i}\\ ~{}~{}\max\limits_{j}~{}~{}\text{d}(v_{i},v_{j}),&\text{otherwise}\end{cases}$ (4)
where $\text{d}(v_{i},v_{j})$ denotes the Euclidean distance. Density peak tokens with high $\rho_{i}\times\delta_{i}$ serve as anchor tokens, indicating that they are surrounded by neighbors with lower local density while remaining relatively distant from other high-density tokens. This selection ensures that anchor tokens are both representative and distinctive. Next, for each anchor frame, Similarity-based Assignment assigns each token in the segment to the nearest anchor token using cosine similarity. Finally, we apply Anchor-centric Aggregation to merge the assigned tokens into their respective anchors, preserving key visual structures through representative tokens. For an anchor $a$ and its associated tokens $[b_{1},\cdots,b_{n}]$, the updated $a^{*}$ is computed as:
$a^{*}=\alpha a+\frac{1-\alpha}{n}\sum_{i=1}^{n}b_{i},$ (5)
where $\alpha$ controls the balance between the anchor token and its associated tokens. In Figure 4, we compare our DTM with Cluster-based Token Merging. In LLMs, RoPE [32] encodes relative positional relationships between tokens, making positional information essential for maintaining the spatiotemporal structure of video tokens. While prior methods [13] rely on Cluster-based Token Merging, they discard positional information, causing RoPE to struggle with encoding spatiotemporal structure. In contrast, our DTM maintains the positional information of merged tokens, enhancing visual context understanding.
@@
Specifically, we extract the [CLS] attention score from the pretrained SigLIP head, $\mathbf{A}\in\mathbb{R}^{H\times W}$, where $H$ and $W$ are the spatial dimensions of frame tokens. Since Video LLMs incorporate pooling (see Figure 2), we apply the same operation to $\mathbf{A}$, resulting in $\bar{\mathbf{A}}\in\mathbb{R}^{\bar{H}\times\bar{W}}$, where $\bar{H}$ and $\bar{W}$ are the pooled spatial dimensions. Finally, for each frame, we select the top $(1-d)rN$ tokens with the highest attention scores to preserve critical visual details.